A solid metal sphere with radius 0.740 m carries a net charge of 0.600 nC.

a) Find the magnitude of the electric field at a point 0.100 m outside the surface of the sphere.

b) Find the magnitude of the electric field at a point inside the sphere, 0.100 m below the surface.
Answer #1

a) E = Kq/r2

E = 540N/C

b) electric field inside is zero
Answer #3

(a)

E=kq/r2=9x109 x 0.6x10-9/0.842=7.65N/C

(b)

E=0